Status Report

NASA Space Station On-Orbit Status 16 Aug 2003

By SpaceRef Editor
August 16, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 16 Aug 2003

All ISS systems continue to function nominally, except as noted previously or below.  Day 112 of ISS residency for the Expedition 7 crew.  First crew rest day of this weekend.

As regularly on Saturdays, CDR Yuri Malenchenko and FE/SO Ed Lu performed the weekly 3-hr. station cleaning.  [This includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]
Malenchenko attended to the daily routine maintenance of the SM SOZh life support system (including ASU toilet facilities), while Ed Lu prepared the daily IMS delta file for automatic export/import to update the database.  Yuri also conducted the weekly inspection of the BRPK-1 air/water condensate separator.

Both crewmembers completed their daily physical exercise program on TVIS treadmill, RED resistive expander and VELO cycle with load trainer.

At 9:10am EDT, the crew conducted the weekly planning conference with the ground, discussing next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners), via S-band/audio.  

Ed and Yuri were thanked for doing a superb job yesterday with the lengthy EMU troubleshooting procedure: “You made us look good…as though we had actually trained you for it here on the ground”.  [Although the troubleshooting didn’t solve the problem, the test did help the ground to home in on a clear signature and rule out a number of component failures. At this point it seems very likely that the suit’s gas trap filter has clogged significantly (which has occurred before, early in the Program’s history). The filter is relatively easy to exchange, but the challenge will be to determine what caused it, so that the new one does not clog also.  Leading suspects are (1) a lubricant from within the EMU, (2) corrosion from the Airlock heat exchanger loop, and/or (3) microbial growth from within the EMU tanks, LCVGs (liquid cooled ventilation garments) or the umbilicals/Airlock Heat Exchanger. Excellent video coverage provided by the crew confirms that the water separator is working correctly.  Also in works are plans to test the other spacesuit, EMU #3005, in an attempt to clear it and the umbilicals of having any related symptoms.] 

For nine hours, starting at 10:00am EDT, JSC/MCC-H is experiencing an outage of its computer firewall due to scheduled maintenance.  [Disrupted during this time are connectivity to other NASA Centers, Russia, other international partners, contractors outside the Clear Lake area, USA and the public Internet.  In addition, the JSC PPTP (point-to-point tunneling protocol) and institutional file and email services are interrupted.  Onsite access to institutional file and email services, inside the firewall, are not affected.]

Weekly Science Update (Expedition Seven — 15th):
The crew was thanked by Lead Increment Scientist Vic Cooley for locating the missing MSG W206 parallel-port cable and for helping to ascertain the ability of the MAMS HiRAP sensor in ER1 to pick up accelerations that might affect experiments in the MSG.  The PFMI checkout went well; all the commands were received and executed as expected.  POIC (Payload Operations & Integration Center) is ready for the first two PFMI sample runs in the coming week which will be looking at pore formation and movement during re-melting as well as possibly some melting kinetics. PFMI will also study interface pattern formation, porosity development and movement within the dendritic array and, possibly, some transient effects on microstructure during re-solidification.

GASMAP:   The 30-day health check for August was completed successfully.  All systems are in great shape and the level of vacuum is exceptional.  Will be done again next month.
Human Research Facility/Workstation (HRF WS):   Continuing.

Ultrasound (USND):  Next Ultrasound session is planned for mid-September.

Pore Formation and Mobility Investigation (PFMI):  Ground is looking forward to beginning science operations again next week.

Space Acceleration Measurement System (SAMS):  SAMS has been offline since 7/31 (3:21pm EDT) and a troubleshooting plan has been developed.  Thanks to the crew for completing the SAMS troubleshooting steps this week which have isolated the problem to the ICU (interface control unit) laptop.  Now looking for a laptop replacement.

Microgravity Acceleration Measurement System (MAMS):  MAMS instrument was nominal throughout the week.  The HiRAP accelerometer has been activated to make vibratory measurements until SAMS is restored to nominal operations. 

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  PCG-STES has survived the RIC (rack interface controller) reboot following the power-down event two weeks ago and is nominal.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):   Current test matrix is complete.

Materials ISS Experiment (MISSE):  In progress.  Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems (CBOSS):  CBOSS-FDI investigation will be scheduled sometime after 12P docking.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Activities are completed for Increment 7.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Activities have been concluded for this Expedition.

Educational Payload Operations (EPO):  The ground thanked the crew for participating in this week’s EVA/IVA tools demonstration.  Excellent feedback on the activity was received.  Everyone is looking forward to future educational demonstrations.

Crew Earth Observation (CEO):   This week’s contribution to Earth Observatory will use the ISS/CEO image of the Mamore River, illustrating the dynamics of its meanders.  CEO team is pleased to report that ISS acquired a splendid view of another of the crew’s favored targets,- The Valley of the Kings, across the Nile River from Luxor.  CEO also got a near-nadir view of the Lewis and Clark Site: Mouth of the Bad River.  While the crew still has some work to do on their technique, a number interesting cities are recognizable, including Manila, Rome, Naples, Cairo, and especially Jidda-Mecca.  The ground enjoyed CEO’s great shots of familiar Texas cities and a beautiful moonrise sequence.  Excellent use was also made of a context view of southwestern Ireland before switching to the longer lens.  This technique is very helpful to cataloging.
Today’s CEO targets (optional), limited in the current XPOP attitude by flight rule constraints on the use of the Lab nadir/science window, and including the targets of the Lewis & Clark 200-year memorial locations, were Howland Island, Central Pacific (coral reef site on the Equator.  Looking a touch right), Nihoa (coral reef site WNW of the big island of Hawaii.  Looking a touch left), Tel Aviv, Israel (pointing a touch left), Amman, Jordan (pointing a touch right), Statue of Zeus at Olympia (looking a touch right on the west coast of the Pelopponese peninsula), Urumqui, China (nadir pass), Smog over the NE USA (Dynamic event.  Major plume exiting into the Atlantic.  Shooting obliques left and right of this ascending track [which followed the coastline just offshore], Temple of Artemis at Ephesus (nadir pass), Mausoleum at Halicarnassus (nadir pass), and Fort Mandan,  ND (LEWIS & CLARK SITE: Nadir pass, on the Missouri River).

See also the website “Space Station Challenge” at

ISS Orbit  (as of this morning, 6:21am EDT [= epoch]):

  • Mean altitude — 383.9 km
  • Apogee — 389.3 km
  • Perigee  378.6 km
  • Period — 92.23 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0007892
  • Orbits per 24-hr. day — 15.61
  • Mean altitude loss in last 24 hours — 75 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 27041
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.